OpenStudy (anonymous):

What is the volume of a cone with a height of 6 centimeters and a radius of 2 centimeters (use 3.14 for )?

OpenStudy (wolf1728):

Cone Volume = (π * r² * h) / 3 Cone Volume = (3.14 * 2*2*6) / 3 That should get you off to a good start

OpenStudy (anonymous):

thank you so much!